Gra 0/201 8 11:59 PM ?55/10003/28201 80807 PM Periocio Table Mup apling Learning A 0.246-kg toy car is powered by one D cell (1.50 V) connected an effective energy conversion efficiency or 41.6% meaning that 4 1.6% of the electric the motor is converted into translational kinetic energy. After 7.88 seconds, the car, which is reaches a speed of 2.35 m/s. What is the average current supplied to the cars motor? to a small DC motor. The car has applied to initially at rest Number O Previous Check Answer Next
Answer #1

Accoridng to the given problem,

[Given]Final KE = 0.5*m*v2 = 0.5*0.246*2.352 J = 0.68 J

So the energy supplied by the battery must be 3.00J / 0.416 = 1.633J

1.633 J is an average power of 1.633 W for 1 second.

Over 7.88 s the average power is therefore 1.633 Ws / 7.88s = 0.2072 W

And Using the power formula,

P = V * I

so the average current is I = P/V = 0.2072W / 1.5V = 0.138 A

I hope you understood the problem, If yes rate me!! or else comment for a better solution.